OpenStudy (anonymous):

5x – 9y = –16 2x + 6y = –16 Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system

OpenStudy (anonymous):

You need one of the variable in the equations to be equal (or a negative version of the other) So that when you add or subtract one equation from the other, that variable is eliminated.

OpenStudy (anonymous):

Can You SHow Me ?? I Have A Big Problem With Trying To Figure OUtt How To Work Outt Substitution & Elimanation

OpenStudy (anonymous):

So equations: (1) 5x-9y=-16 (2) 2x+6y=-16 We can make them both have 10 as their x coefficients (1) 2*(5x-9y=-16) (2) 5*(2x+6y=-16) simplify (1)10x - 18y = - 32 (2)10x + 30y =-80 Equation 1 minus equation 2: -48y = 48 y=-1

OpenStudy (anonymous):

Oh Ok =]

OpenStudy (anonymous):

Thanxx Alot ! =]